In a certain box of gloves, 12 pairs are size 7 and 24 pairs are size 6. If all the gloves in the box are either size 6 or 7, what percent of the gloves in the box are size 6?
Answer by richwmiller(17219)

24/(12+24)=24/36=2/3= 66.67%